OpenStudy (anonymous):

Numbers p and q are positive integers. Which statements are always true? Choose all answers that are correct. A. p + q is a positive integer B. p – q is a negative integer C. p • q is a positive integer D. p/q is a positive integer @Haseeb96 @iGreen @Joel_the_boss @LunyMoony @LazyBoy

OpenStudy (lunymoony):

A, C, D

OpenStudy (lunymoony):

If they are all positive, it is those 3

OpenStudy (haseeb96):

lets we suppose p =2 and q =4 so when we add these two number we get also a positive integer

OpenStudy (haseeb96):

do u got it ?

OpenStudy (haseeb96):

@LunyMoony go give her direct answer

OpenStudy (anonymous):

yup

OpenStudy (haseeb96):

so first answer is ____ ???@Wolfie_GirlRocks2002

OpenStudy (lunymoony):

you mean "dont" ?

OpenStudy (haseeb96):

now second choice p-q is a negative integer 2-4 = -2 but when p=9 and q=6

OpenStudy (anonymous):

I know that you said P=2 and Q=4 add them together and you get a positive so it would be A

OpenStudy (haseeb96):

yes 9-6=3 so 3 is not negative integer so B is not our answer because it does not satisfy

OpenStudy (anonymous):

I don't get that one

OpenStudy (haseeb96):

when p=2 and q =3 p.q is a positive integer 2.3=6 6 is also positive integer

OpenStudy (haseeb96):

so C is also our answer

OpenStudy (haseeb96):

wait i will tell u about B choice in the end

OpenStudy (anonymous):

what about d?

OpenStudy (haseeb96):

p/q is also a positive integer p=2 and q=4 2/4=1/2 =0.5 0.5 is also a positive integer so D is also our answer

OpenStudy (anonymous):

so we have A,C, and D

OpenStudy (haseeb96):

look here now second choice ''p-q is a negative integer'' p=2 and q=4 2-4 = -2 -2 is a negative number but

OpenStudy (anonymous):

it is A an C only

OpenStudy (haseeb96):

when we have these type of number when p=9 and q =6 p-q is a negative number 9-6=3 but here 3 is a positive number not negative so we can say that B is not our answer

OpenStudy (haseeb96):

yes D is not our answer sorry i forgot @surjithayer thanks dear

OpenStudy (anonymous):

if p=4,q=3 4+3=7 (positive integer) 4-3=1(not a negative integer) 4*3=12 (positive integer) 4/3 (not an integer.)

OpenStudy (anonymous):

thanks for catching that

OpenStudy (anonymous):

0.5 is not an integer.

OpenStudy (anonymous):

yw